OpenStudy (anonymous):

Find the resultant displacement of a bear searching for berries on the mountain. The bear heads 55.0º north of west for 15.0 m; then it turns and heads to the west for another 7.00 m.

OpenStudy (badhi):

 find length of OQ by using cosine rule. ie., $$OQ^2=7^2+15^2+2(7)(15)\cos \alpha$$ where $$\alpha+55^\circ=180^\circ$$